OpenStudy (jimmiejones153):

(06.04) 4x + 1 over 3y2 What is the value of the expression above when x = 2 and y = 3? You must show all work and calculations to receive full credit.

OpenStudy (supersmart1001):

substituting your numerical data into your expression, we get: 4x+13y2=4⋅2+13⋅32=...

OpenStudy (supersmart1001):

hold on let m rewrite that

OpenStudy (jimmiejones153):

ok

OpenStudy (supersmart1001):

4x+1/3y^2=4.2+1/3*3^2

OpenStudy (supersmart1001):

you have to do multiplications first

OpenStudy (jimmiejones153):

ok

OpenStudy (supersmart1001):

first step: 4⋅2=...

OpenStudy (jimmiejones153):

8

OpenStudy (supersmart1001):

ok! then: 1/3⋅9=...?

OpenStudy (jimmiejones153):

3